Please complete thank you
Q3) (based on textbook problem 3.26) Annie consumes three goods out of her income, food (F) shelter (S), and clothing (C). At her current levels of consumption, her marginal rate of substitution of food for clothing is 3 and her marginal rate of substitution of clothing for shelter is 2. a) What is her marginal rate of substitution of food for shelter? (Hint: Use the fact that MRS is a ratio of marginal utilities) b) Do you have enough information to determine her marginal utility of shelter? If so, what is it? If not, why not?
Answer #1

Marginal rate of substitutionFC = MUF/MUC

MRSFC = MUF / MUC = 3

MRSCS = MUC / MUS = 2

A. Her marginal rate of substitution food for shelter

large rac{MU_F}{MU_C}*rac{MU_C}{MU_S}= 3*2

MRS$ MUS

Marginal rate of substitution food for shelter = 6.

B. No we cannot determine the marginal utility of shelter. Here if we were given Price of the commodity along with total income then it was possible to calculate marginal utility of different commodities.

The MES gives us only quantitative information that is ratio of marginal utility. Hence we cannot determine Marginal utility of Shelter.

Please contact if having any query thank you. Kindly rate up if satisfied. Thank you for your generous support